a bottle of medicine must be stored at a temperature between 50'f and 86'f. what range does this correspond to on the celsius scale? c=5/9(f-32)]
Answer by Boreal(15235)

C=(5/9)(50-32)=5/9*18=10 deg C
similarly 86 F= 30 deg C.
0 C====32 F
10C===50F 10 to 18 ratio
20C===68F
30C===86F
40C===104 F
-40C=-40F only place they are the same